1. Find general solutions for the equations: (a) y - 4y - 5y (b) y + 3y + 4y = 0.
Answer #1

yl)-44. 5y = 0 Auxillary ean is m² - 4m-50 =0 m² - 5m + m - 5=0 mln - 5) + (-5) =0. m =1 General son - y m =-5 is not = eie +